Apologies for absence
Minutes of the last meeting.
These were accepted.
We may have to turn racks around in the Forum server room to improve the air flow. We would also like to make the floor grills level with the surrounding floor tiles. A trolley with an integrated monitor panel would be useful. Infrastucture unit will coordinate, but wait for new TP leads so can rewire at the same time.
Switch naming unification will be added to DevProj #33.
We have a number of staff leaving (and some that have already left) that will not be replaced any time soon due to a staff recruitment freeze. It is critical that we retain development effort - so we should review use of central services and adopt as pro-active as possible awareness of what IS are working on - some areas of IS are better than others at communicating work and requirements capture on projects. There is not likely to be any formal review of Units and effort allocation until after the staff recruitment freeze is lifted.
Despite not moving from JCMB imminently due to the air conditioning problems in the Forum server room there will still be a gap to cover for offsite provision.
May need to turn off non-essential machines and equipment in the Forum server room if the air conditioning issues are not resolved and are going to be a frequent occurence.
There should be a policy document on DICE web site.
Support vaguely look after the A/V cabinets in meeting rooms, although they have not been officially handed over to the School yet. There will be a DICE box installed in each with wireless keyboard and mouse. A remote USB connection for the SmartBoard may be useful.
Reports from units.
This was the first critical use of the DChW, up until now if its been broken Works have not been bothered about fixing it quickly - this attitute should change now. In this case both backup systems failed - the backup for the DChW and our own Forum server room backup.
Will decant (probably over Christmas) to either Forum or FH.
Toby will circulate timetable for installing the new certificate. Services using the old apache component may need to be manually restarted to pick up new X509 certificates.
There are no remaining web services using kx509 authentication, so the authportal will be decommissioned on 11/09. The only remaining service that uses kx509 is openvpn.
Server reboots for 5.2 upgrade will not be automatic - individual Units will need to schedule.
PXE works but new install CD will be needed.
Reference to "network" component should be to "routing" component. It will need to be fixed to allow static route to be added.
All 170-odd tapes will need to be wiped before bringing this into service. May need some tapes to be physically relocated for backups of machines remaining at JCMB. Need written down official disaster recovery procedures noting unit responsibilities.
Resolution rate for older tickets in non-support queues has dropped, units may wish to check on progress.
The following points were agreed/discussed:
The following points emerged from the discussion:
The next meeting will be on the 24th of September and chaired by Tim Colles. Time and location to be confirmed.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh